# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v16

@@ Abstract
As a rigorous statistical approach, statistical Taylor expansion extends the conventional Taylor expansion by replacing precise input variables with random variables of known distributions to compute the means and standard deviations of the results. Statistical Taylor expansion tracks the propagation of the input uncertainties through intermediate steps, causing the variables in intermediate analytic expressions to become interdependent, while the final analytic result becomes path independent. This fundamentally distinguishes it from common approaches in applied mathematics that optimize computational path for each calculation. In essence, statistical Taylor expansion may standardize numerical computations for analytic expressions. Its statistical nature enables reliable testing of results when the sample size is sufficiently large, as well as includes sample count in result qualification. This study also introduces an implementation of statistical Taylor expansion termed variance arithmetic and presents corresponding test results across a wide range of mathematical applications.
Another important conclusion of this study is that numerical errors in library functions can significantly affect results. For instance, periodic numerical errors in trigonometric library functions may resonate with periodic signals, producing large numerical errors in the outcomes.
@@ 1 Introduction
1.1 Measurement Uncertainty
Except for the most basic counting, scientific and engineering measurements never yield completely precise results [1][2]. In such measurements, the uncertainty of a quantity $x$ is typically expressed either by the sample deviation $\delta x$ or by the uncertainty range $\Delta x$ [1][2].
  • If $\delta x=0$ or $\Delta x=0$, $x$ is a precise value.
  • Otherwise, $x$ is an imprecise value.
$P(x)\equiv\delta x/|x|$ is defined as the statistical precision (hereafter referred to as precision) of the measurement, where $x$ denotes the value, and $\delta x$ represents the uncertainty deviation. A larger precision indicates a coarser measurement whereas a smaller precision indicates a finer measurement. The precision of measured values can range from order-of-magnitude estimates to $10^{-2}$ to $10^{-4}$ in common measurements, and to $10^{-15}$ in state-of-the-art determinations of fundamental physical constants [3].
This study focuses on determining the result value and uncertainty of a general analytic expression with imprecise input values.
1.2 Extension to Existing Statistics
Rather than focusing on the uncertainty distribution of the result when applying a function to a random variable with a known uncertainty distribution [4], statistical Taylor expansion provides the mean and variance of the result for general analytic expressions.
@@ 1.2 Extension to Existing Statistics
  • The traditional variance-covariance framework accounts only for linear interactions between random variables through an analytic function [4][5], whereas statistical Taylor expansion extends this framework to include higher-order interactions as expressed in Formula (2.3) in this paper. As shown in Formula (7.9), (7.17), and (7.18), such extension is important even for linear algebra.
1.3 Problem of the Related Numerical Arithmetics
Variance arithmetic implements statistical Taylor expansion and surpasses all existing related numerical arithmetic methods.
1.3.1 Conventional Floating-Point Arithmetic
Conventional floating-point arithmetic [6][7][8] treats every bit of its significand to be valid at all times, and continuously either generates artificial information or erases useful information to maintain this assumption during the normalization process [9]. For example, the following calculation is performed exactly in integer format:
$64919224\times 205117922-159018721\times 83739041=\\ 13316075197586562-13316075197586561=1;$ (1.1)
If Formula (1.1) is carried out using conventional 64-bit floating-point arithmetic:
$64919224\times 205117922-159018721\times 83739041=\\ 64919224.000000000\times 205117922.000000000-159018721.000000000\times 83739041.000000000=\\ 13316075197586562.-13316075197586560.=2.=2.0000000000000000;$ (1.2)
@@ 1.3.1 Conventional Floating-Point Arithmetic
    The normalization of input values may artificially pad zero bits in the least significant positions until every bit in the 53-bit significand is used, for example, converting $64919224$ to $64919224.000000000$.
@@
    The results of addition or multiplication may exceed the bit range of the 53-bit significand, requiring rounding off from the least significant bits to fit within the 53-bit significand. This process can generate rounding errors, for example, converting $13316075197586561$ to $13316075197586560$, which produces a rounding error of $1$ in the least significant bit.
@@
    The results of subtraction or division can cancel values from higher significant bits, exposing rounding errors in the lower significant bits. For example, 13316075197586562.13316075197586560.=213316075197586562.-13316075197586560.=2.
@@
    The normalization of subtraction or division results can amplify rounding errors toward the most significant bit by padding zeros, such as converting $2.$ to $2.0000000000000000$.
Formula (1.2) demonstrates a classic case of catastrophic cancellation [10][11].
Because rounding errors from lower digits can propagate to higher digits, the $10^{-7}$ significance of the 32-bit IEEE floating-point format [6][7][8] is usually insufficient for calculations involving input data with a precision of $10^{-2}$ to $10^{-4}$. For more complex calculations, even the $10^{-16}$ significance of the 64-bit IEEE floating-point format [6][7][8] may be sufficient for inputs with $10^{-2}$ to $10^{-4}$ precision. This represents a fundamental controversy of conventional floating-point arithmetic.
Because rounding error is path dependent, a major objective of conventional numerical methods is to identify optimal computation strategies that minimize rounding errors, such as Gaussian elimination [11][12][13], even though all alternative paths of root finding methods in linear algebra are mathematically equivalent [14].
Self-censoring rules have been developed to limit such rounding error propagation [12][13], for example, by avoiding the subtraction of results from large multiplications, as shown in Formula (1.2). However, these rules are neither enforceable nor easily adoptable, and they are even more difficult to quantify. To date, research on rounding error propagation has focused primarily on linear calculations [10][11][15], or on special cases [12][16][17], whereas in practice, rounding errors often manifest as pervasive and mysterious numerical instabilities [18].
The forward rounding error study [11] compares (1) the result containing rounding error and (2) the ideal result without rounding error, for example, by comparing a result obtained using 32-bit IEEE floating-point arithmetic with the corresponding result computed using 64-bit IEEE floating-point arithmetic [19]. The most recent study of this type presents an extremely optimistic view of numerical library errors, reporting them as fractions of the least significant bit of the floating-point significand [19]. However, such optimism contradicts the statistical tests on numerical library functions presented in this paper.
The backward rounding error study [10][11][15] estimates only the result uncertainty caused by rounding errors, thereby overlooking the bias that rounding errors introduce into the result value. This analysis is typically limited to very small uncertainties because it relies on perturbation theory and it is tailored to each specific algorithm [10][11][15].
In contrast, variance arithmetic traces rounding error directly as part of uncertainty. Statistical Taylor expansion applies generally to any analytic function, providing both result mean and deviation, and accommodates input uncertainties of any magnitude. By demonstrating that the analytic result should be path independent, statistical Taylor expansion fundamentally challenges the conventional methodology of seeking an optimal execution strategy for a given analytic expression.
1.3.2 Interval Arithmetic
Interval arithmetic [13][20][21][22][23][24] is currently a standard method for tracking computational uncertainty. Its objective is to ensure that a value remains strictly bounded within its bounding range throughout the computation.
However, the bounding range in interval arithmetic is not compatible with the approach commonly used in scientific and engineering measurements, which instead characterizes uncertainty in terms of the statistical mean and deviation [1][2]  [^2] There is one attempt to connect intervals in interval arithmetic to confidence interval or the equivalent so-called p-box in statistics [25]. Because this attempt seems to rely heavily on 1) specific properties of the uncertainty distribution within the interval and/or 2) specific properties of the functions upon which the interval arithmetic is used, this attempt does not seem to be generic. If probability model is introduced to interval arithmetic to allow tiny bounding leakage, the bounding range is much less than the corresponding pure bounding range [15]. Anyway, these attempts seem to be outside the main course of interval arithmetic. .
Interval arithmetic represents only the worst-case scenario of uncertainty propagation. For example, in addition, it assumes that the two input variables are perfectly positively correlated [26], thereby producing the widest possible bounding range. In contrast, if the variables were perfectly negatively correlated, the bounding range after addition would decrease [27]  [^3] Such case is called the best case in random interval arithmetic. The vast overestimation of bounding ranges in these two worst cases prompts the development of affine arithmetic [26][28], which traces error sources using a first-order model. Being expensive in execution and depending on approximate modeling even for such basic operations as multiplication and division, affine arithmetic has not been widely used. In another approach, random interval arithmetic [27] randomly chooses between the best-case and the worst-case intervals, so that it can no longer guarantee bounding without leakage. Anyway, these attempts seem to be outside the main course of interval arithmetic. . This worst-case assumption can lead to order-of-magnitude over-estimations [29].
The results of interval arithmetic can depend strongly on the specific algebraic form of an analytic function $f(x)$, a phenomenon known as the dependency problem. This issue is amplified in interval arithmetic [23] but also exists in conventional floating-point arithmetic [12].
Furthermore, interval arithmetic lacks a mechanism to reject invalid calculations, even though every mathematical operation has a valid input range. For example, it produces branched results for $1/(x\pm\Delta x)$ or $\sqrt{x\pm\Delta x}$ when $0\in[x-\Delta x,x+\Delta x]$, whereas a context-sensitive uncertainty bearing arithmetic should reject such calculation naturally.
In contrast, variance arithmetic specifies each value by its mean and deviation. It does not suffer from the dependency problem. Its statistical framework naturally rejects certain input intervals on mathematical grounds, such as inversion and square root operations when the statistical bounding range includes zero.
1.3.3 Statistical Propagation of Uncertainty
Statistical propagation of uncertainty treats each imprecise value as a random variable and seeks correlations among such random variables to calculate resulting mean and deviation [30].
In contrast, statistical Taylor expansion interprets the uncertainty of an imprecise value as a limitation in obtaining its precise value. This assumption aligns with most error analyses in the literature [1][2]. Its statistical foundation is the uncorrelated uncertainty assumption [29]: any two imprecise values do not correlate in their uncertainties.
Statistical propagation of uncertainty appears to have applied an incorrect statistical context to uncertainty. For example, a time series is a random variable whereas each imprecise value within the time series is merely an imprecise measurement. The uncorrelated uncertainty assumption can hold for each imprecise value in a time series, even though the time series can highly correlate to another time series at their signal level, or even correlated to itself as a periodic time series. If the uncertainties of some imprecise values are correlated, these imprecise values contain systematic errors among them [1]. Systematic errors should be treated as unwanted signals but not as normal uncertainties [1]. In short, uncertainty should not be treated as signal.
1.3.4 Significance Arithmetic
Significance arithmetic [31] seeks to track the number of reliable bits in an imprecise value throughout a calculation. In its early implementations [32][33], significance arithmetic was based on simple operational rules applied to reliable bit counts rather than on formal statistical methods. In these approaches, the reliable bit count is treated as an integer, even though in practice it can take a fractional value [34]. This constraint can produce an artificial step-wise reduction in significance. The implementation of significance arithmetic in Mathematica [34] employs a linear error model consistent with the first-order approximation of interval arithmetic [13][22][23].
One limitation of significance arithmetic is its inability to specify uncertainty accurately [29]. For example, if the least significant bit of the significand is used to represent uncertainty, the result precision can be very coarse, as in $1\pm 10^{-3}=1024\times 2^{-10}$ [29]. Introducing a limited number of bits to represent uncertainty does not fully resolve this issue. Consequently, various attempts to develop floating-point arithmetic without normalization [33] have not been widely adopted, and conventional floating-point arithmetic continues to dominate the numerical world. For this reason, variance arithmetic has abandoned the significance arithmetic principles of its predecessor [29].
1.3.5 Stochastic Arithmetic
Stochastic arithmetic [35][36] randomizes the least significant bits of each input floating-point value, repeats the same calculation multiple times, and then applies statistical analysis to identify the invariant digits among the results as significant digits. However, this approach can be computationally expensive as the number of repetitions required for each input depends on the algorithm and can increase substantially when the algorithm contains branching operations.
In contrast, statistical Taylor expansion provides a direct characterization of the result’s mean and deviation without sampling.
1.4 An Overview of This Paper
This paper presents the theory of statistical Taylor expansion, its implementation as variance arithmetic, and the corresponding validation tests. Section 1 compares statistical Taylor expansion and variance arithmetic with other established uncertainty bearing arithmetic. Section 2 develops the theoretical foundation of statistical Taylor expansion. Section 3 describes variance arithmetic as an implementation of statistical Taylor expansion. Section 4 outlines the standards and methodologies used to validate variance arithmetic. Section 5 illustrates variance arithmetic in computing polynomial. Section 6 evaluates variance arithmetic on common mathematical library functions. Section 7 applies variance arithmetic to adjugate matrix and matrix inversion. Section 8 demonstrates its application to time-series data. Section 9 examines the impact of numerical library errors and shows that these errors can be significant. Section 10 applies variance arithmetic to regression analysis. Section 11 concludes with a summary and discussion.
@@ 2 Statistical Taylor Expansion
2.1 The Uncorrelated Uncertainty Assumption
Figure 1: Effect of noise on bit values of a measured value. The triangular wave signal and the added white noise are shown at top using the thin black line and the grey area, respectively. The values are measured by a theoretical 4-bit Digital-to-Analog Converter in ideal condition, assuming LSB is the 0th bit. The measured 3rd and 2nd bits without the added noise are shown using thin black lines, while the mean values of the measured 3rd and 2nd bits with the added noise are shown using thin grey lines. This figure is a reproduction of Figure 1 in [29]
Figure 2: $\gamma$ vs $P$ for different $\gamma_{P}$ in Formula (2.4). If $\gamma_{P}$ is the maximal allowed correlation on the uncertainty level for uncorrelated uncertainty assumption to hold, then $\gamma$ is the maximal allowed correlation for two input to hold. For a given $\gamma_{P}$, $\gamma$ increases with finer $P$. This figure is a reproduction of Figure 2 in [29].
The uncorrelated uncertainty assumption [29] states that the uncertainties of any two input imprecise values are uncorrelated. This assumption is satisfied when there is no systematic errors in the inputs [1] and is consistent with standard methods for processing experimental data [1][2]. The uncorrelated uncertainty assumption permits the signals themselves to be highly correlated in the following statistical test [29].
Let $X$, $Y$, and $Z$ denote three mutually independent random variables [4] with variance $V(X)$, $V(Y)$ and $V(Z)$, respectively. Let $\alpha$ denote a constant. Let $Cov()$ denote the covariance function. Let $\gamma$ denote the correlation between $(X+Y)$ and $(\alpha X+Z)$. And let:
$\eta_{1}^{2}\equiv\frac{V(Y)}{V(X)};\;\;\;\eta_{2}^{2}\equiv\frac{V(Z)}{V(\alpha X)}=\frac{V(Z)}{\alpha^{2}V(X)};$ (2.1)
$\gamma=\frac{Cov(X+Y,\alpha X+Z)}{\sqrt{V(X+Y)}\sqrt{V(\alpha X+Z)}}=\frac{\alpha/|\alpha|}{\sqrt{1+\eta_{1}^{2}}\sqrt{1+\eta_{2}^{2}}}\equiv\frac{\alpha/|\alpha|}{1+\eta^{2}};$ (2.2)
Formula (2.2) gives the correlation $\gamma$ between two random variables, each of which contains a completely uncorrelated part and a completely correlated part $X$, with $\eta$ being the average ratio between these two parts. Formula (2.2) can also be interpreted reversely: if two random variables are correlated by $\gamma$, each of them can be viewed hypothetically as containing a completely uncorrelated part and a completely correlated part, with $\eta$ being the average ratio between these two parts  [^4] The widely used common factor analysis is also based on this principle.. The correlated parts common to different measurements are regarded as signals, which can either be desired or unwanted.
One special application of Formula (2.2) is the correlation between a measured signal and its true signal, in which noise is the uncorrelated part between the two. Figure 1 shows the effect of noise on the most significant two bits of a 4-bit measured signal when $\eta=1/4$. Its top chart shows a triangular waveform between 0 and 16 as a black line, and a white noise between -2 and +2, using the grey area. The measured signal is the sum of the triangle waveform and the noise. The middle chart of Figure 1 shows the values of the 3rd digit of the true signal as a black line, and the mean values of the 3rd bit of the measurement as a grey line. The 3rd bit is affected by the noise during its transition between 0 and 1. For example, when the signal is slightly below 8, only a small positive noise can turn the 3rd digit from 0 to 1. The bottom chart of Figure 1 shows the values of the 2nd digit of the signal and the measurement as a black line and a grey line, respectively. Figure 1 clearly shows that the correlation between the measurement and the true signal is less at the 2nd digit than at the 3rd digit. Quantitatively:
  • The overall measurement is 99.2% correlated to the signal with $\eta=1/8$;
  • The 3rd digit of the measurement is 97.0% correlated to the signal with $\eta=1/4$;
  • The 2nd digit of the measurement is 89.4% correlated to the signal with $\eta=1/2$;
  • The 1st digit of the measurement is 70.7% correlated to the signal with $\eta=1$;
  • The 0th digit of the measurement is 44.7% correlated to the signal with $\eta=2$.
The above conclusion agrees with the common experiences that, below the noise level of measured signals, noises rather than true signals dominate each digit.
Similarly, while the correlated portion between two values has the same value at each bit of the two values, the ratio of the uncorrelated portion to the correlated portion increases by 2-fold for each bit down from MSB of the two values. Quantitatively, let $P$ denote the precision of an imprecise value, and let $\eta_{P}$ denote the ratio of the uncorrelated portion to the correlated portion at level of uncertainty; then $\eta_{P}$ increases with decreased $P$ according to Formula (2.3). According to Formula (2.2), if two significant values are overall correlated with $\gamma$, at the level of uncertainty the correlation between the two values decreases to $\gamma_{P}$ according to Formula (2.4).
$\eta_{P}=\frac{\eta}{P},\;\;\;P<1;$ (2.3)
$\frac{1}{\gamma_{P}}-1=\left(\frac{1}{\gamma}-1\right)\frac{1}{P^{2}},\;\;\;P<1;$ (2.4)
Figure 2 plots the relation of $\gamma$ vs. $P$ for each given $\gamma_{P}$ in Formula (2.4). When $\gamma_{P}$ is less than a predefined maximal threshold (e.g., 2%, 5% or 10%), the two values can be deemed uncorrelated to each other at the level of uncertainty. For each independence standard $\gamma_{P}$, there is a maximal allowed correlation between two values below which the uncorrelated uncertainty assumption of statistical Taylor expansion holds. The maximal allowed correlation is a function of the larger precision of the two values according to Formula (2.4). Figure 2 shows that for two precisely measured values, their correlation $\gamma$ is allowed to be quite high.
@@ 2.2 Distributional Zero and Distributional Pole
$y=x^{c}:\;\;\;\rho(\tilde{y},\mu_{y},\sigma_{y})=c\tilde{y}^{\frac{1}{c}-1}\frac{1}{\sigma}N(\frac{\tilde{y}^{\frac{1}{c}}-\mu}{\sigma});$ (2.6)
@@
  • If $\tilde{y}=\alpha+\beta\tilde{x}$, the resulting distribution is identical to the original distribution, since $\rho(\tilde{y},\mu_{y},\sigma_{y})=\rho(\tilde{y},\alpha+\beta\mu,\beta\sigma)$ [4]. A linear transformation generates neither a distributional zero nor a distributional pole, according to Formula (2.5).
  • Figure 3 shows the probability density function for $(x\pm 1)^{2}$ according to Formula (2.6), which exhibits a distributional pole at $x=0$. The distribution $(0\pm 1)^{2}$ corresponds to the $\chi^{2}$ distribution [1]. At the distributional pole, the probability density function resembles a Delta distribution.
  • Figure 4 illustrates the probability density function for $\sqrt{x\pm 1}$ according to Formula (2.6), which has a distributional zero at $x=0$. At the distributional zero, the probability density function is zero.
In both Figure 3 and 4, $\rho(\tilde{y},\mu_{y},\sigma_{y})$ closely representation resembles $\rho(\tilde{x},\mu,\sigma)$ when the mode of $\rho(\tilde{x},\mu,\sigma)$ lies sufficiently far away from either a distributional pole or a distributional zero, thereby allowing for a generic characterization of the output.

### v17

@@ Abstract
As a rigorous statistical approach, statistical Taylor expansion extends the conventional Taylor expansion by replacing precise input variables with random variables of known distributions and sample counts to compute the mean, the standard deviation, and the reliable factor of each result. It tracks the propagation of the input uncertainties through intermediate steps, so that the final analytic result becomes path independent. Therefore, it differs fundamentally from common approaches in applied mathematics that optimize computational path for each calculation. Statistical Taylor expansion may standardize numerical computations for analytic expressions. This study also introduces the implementation of statistical Taylor expansion termed variance arithmetic and presents corresponding test results across a wide range of mathematical applications.
Another important conclusion of this study is that numerical errors in library functions can significantly affect results. It is desirable that each value from library functions be accomplished by an uncertainty deviation.
@@ 1 Introduction
Let $x$ and $\delta x$ be a value and its uncertainty deviation, respectively. As an input, $x$ and $\delta x$ are the mean and standard deviation of a measurement [1][2]. If $\delta x=0$, $x$ is a precise value, otherwise, the pair specifies an imprecise value $x\pm\delta x$.
Let $P(x)\equiv\delta x/|x|$ be the statistical precision (hereafter referred to as precision) of $x\pm\delta x$. A smaller $P(x)$ indicate a higher measurement quality of $x\pm\delta x$, and when $P(x)\ll 1$, it is unclear whether $\delta x$ is too poor to measure $x$, or it is a measurement of $x=0$.
Statistical Taylor expansion determines the result $f\pm\delta f$ of a general analytic expression $f(x,\dots)$, and the reliability $\alpha\in[0,1]$ of $f\pm\delta f$ based on the inputs $x\pm\delta x,\dots$ and their corresponding sample counts.
@@
  • The traditional variance-covariance framework accounts only for linear interactions between random variables through an analytic function [3][4][5], whereas statistical Taylor expansion extends this framework to include higher-order interactions as expressed in Formula (2.3) in this paper.
  • Calculating the reliability of $f\pm\delta f$ seems completely new.
Statistical Taylor expansion is superior to all existing numerical arithmetic.
Conventional floating-point arithmetic [6][7][8] can only calculate the result $f$ which contains unknown amount of rounding error [9][10][11]. As a consequence, a floating-point representation with $10^{-16}$ resolution may not be good enough for inputs with $10^{-2}$ to $10^{-6}$ precision. While statistical Taylor expansion can track rounding errors.
The bounding range in interval arithmetic [12][13][14][15][16] is inconsistent with the statistical nature of an $x\pm\delta x$ pair and it tends to over-estimate the result uncertainty due to its worst-case assumption [17]. On the other hand, statistical Taylor expansion is precise statistically.
Both conventional floating-point arithmetic and interval arithmetic depend strongly on the specific algebraic form of an analytic function, a phenomenon known as the dependency problem [10][11][12][14], which makes the numerical calculations of analytic problems sometimes more like an art than science. In contrast, statistical Taylor expansion is path independent.
To be rigorous in mathematics and statistics, statistical Taylor expansion abandons the significance arithmetic nature of its processor [17].
As a statistical sampling process, stochastic arithmetic [18][19] is computationally expensive, while statistical Taylor expansion provides a direct characterization without sampling.
In this paper:
@@
    Section 1 compares briefly statistical Taylor expansion with all other numerical arithmetic.
@@
    Section 2 develops the theoretical foundation of statistical Taylor expansion.
@@
    Section 3 describes variance arithmetic as a numerical implementation of statistical Taylor expansion.
@@
    Section 4 lays out standard to validate variance arithmetic.
  5. 5.
    Section 5 illustrates variance arithmetic in computing polynomial, demonstrating its capability in tracing floating-point rounding errors, and its continuity in parameter space.
  6. 6.
    Section 6 applies variance arithmetic to adjugate matrix and matrix inversion, showing that the uncertainty calculation for a linear calculation is no longer linear. It also distinguishes between a distribution test and a value test.
  7. 7.
    Section 7 evaluates variance arithmetic on common mathematical library functions. It also shows that the error deviation should be close to $1$ except near a distributional pole.
  8. 8.
    Section 8 reveals the accumulation of numerical errors by reusing an input multiple time in a moving-window progressive algorithm.
  9. 9.
    Section 9 showcases variance arithmetic in catching catastrophic cancellation in a regression algorithm.
  10. 10.
    Section 10 examines the impact of numerical library errors and shows that these errors can be significant.
  11. 11.
    Section 11 concludes with a summary and a discussion.
@@ 2 Statistical Taylor Expansion
2.1 The Uncorrelated Uncertainty Condition
When there is no systematic error [1][2] among the inputs, the uncertainties of the inputs are uncorrelated to each other, even though the inputs may have significant correlations among them. This condition can be characterized quantitatively in statistics as the uncorrelated uncertainty condition [17].
@@ 2.2 Distributional Zero and Distributional Pole
  • A distributional pole occurs when $f^{(1)}_{x}=0\rightarrow\rho(\tilde{y},\mu_{y},\sigma_{y})=\infty$. For example, $(\mu\pm 1)^{2}$ has a distributional pole at $\tilde{x}=0$, as shown in Figure 1.
  • A distributional zero occurs when $f^{(1)}_{x}=\infty\rightarrow\rho(\tilde{y},\mu_{y},\sigma_{y})=0$. For example $\sqrt{\mu\pm 1}$ has a distributional zero at $\tilde{x}=0$, as shown in Figure 2.
In both Figure 1 and 2, $\rho(\tilde{y},\mu_{y},\sigma_{y})$ closely representation resembles $\rho(\tilde{x},\mu,\sigma)$ when the mode of $\rho(\tilde{x},\mu,\sigma)$ lies sufficiently far away from either a distributional pole or a distributional zero, such as $(5\pm 1)^{2}$ in Figure 1 and $\sqrt{5\pm 1}$ in Figure 2, thereby allowing for characterization of the output $f(x)$ using mean $\overline{f(x)}$ and deviation $\delta f(x)$ only.